# Stabilizing black-box model selection with the inflated argmax

### v1

@@ 3.1 Alternative selection rules
Stability selection.
In the context of variable selection, the stability selection method (Meinshausen and Bühlmann,, 2010) selects a single model based on whether each covariate is selected in at least $\tau$ proportion of bags. Concretely, given a vector of weights $\hat{w}$ indexed by $M^{+}=2^{[d]}$, define the marginal variable probabilities $\hat{\Pi}_{j}=\sum_{\begin{subarray}{c}m\subseteq[d]:j\in m\end{subarray}}\hat{w}_{m},$ for $j=1,\ldots,d.$ Stability selection returns
$\text{ip}_{\tau}(\hat{w})$ =$=\left\{\hat{m}_{\tau}\right\},\text{ where }\hat{m}_{\tau}=\Big{\{}j\in[d]:\hat{\Pi}_{j}\geq\tau\Big{\}},$
for some use-specified threshold $\tau$. Similar to the standard argmax (without ties), $\text{ip}_{\tau}(\hat{w})$ only returns a single model, which can be unstable if some marginal probabilities $\hat{\Pi}_{j}$ are near the threshold $\tau$. It also fails to adapt to relationships among variables (§ 4.1).
4 Examples: variable selection in linear models
Consider the problem of variable selection in linear regression (a specific type of model selection); this is a classic problem in bioinformatics (Saeys et al.,, 2007), environmental studies (Effrosynidis and Arampatzis,, 2021), psychology (Vrieze,, 2012), and more.
In this setting, the data is modelled as the form
$\hat{m}(\mathcal{D})=\mathbbm{1}\{|\hat{\beta}|>0\}$ (8)
where $\hat{\beta}$ implicitly depends on the dataset $\mathcal{D}$. This is an instance of the setting in (2). In this context, the LASSO provides unstable model selections, particularly in settings with nontrivial correlations among covariates (Meinshausen and Bühlmann,, 2010). Moreover, since many variable selection algorithms, including the LASSO, only return a single set of selected covariates, these methods do not directly provide an approach for assessing uncertainty in the selected model.
4.1 Linear regression with correlated covariates
@@
4.2 Sparse identification of nonlinear dynamical systems (SINDy)
One application of sparse regression, highlighted in our experiments, is the discovery of governing equations of a system from observations. Specifically, we seek to learn the governing equations of a system of the form
$\frac{d}{dt}u(t)=f(u(t)),$ (13)
where $u(t)=\begin{bmatrix}u^{(1)}(t),u^{(2)}(t),\cdots,u^{(d)}(t)\end{bmatrix}^{\top}$ is a $d$-dimensional state vector containing quantities of interest at time $t$, and $f$ is the function describing how the derivative of $u(t)$ evolves with time. The Sparse Identification of Nonlinear Dynamical Systems (SINDy, Brunton et al., (2016)) algorithm represents a differential equation of the form (13) as a weighted sum of terms in a library of functional forms, formulates the problem as an instance of linear regression (6), and solves (7) or a related method to “discover” the differential equations $f$ from data. Specifically, $y_{i}$ in (6) may correspond to estimates of $\frac{d}{dt}u(t_{i})$ at the $i$-th observation time, and $x_{i}$ may correspond to a vector of functionals of $u(t_{i})$, with elements such as $(u^{(1)}(t_{i}))^{3}$ or $u^{(1)}(t_{i})u^{(3)}(t_{i})$.
Bagging Ensemble SINDy (E-SINDy, Fasel et al., (2022)) creates an ensemble of model fits, and with this ensemble, the final model is chosen. This approach corresponds to applying stability selection (Meinshausen and Bühlmann,, 2010; Shah and Samworth,, 2013) to SINDy, using sequentially thresholded ridge regression (STRidge, Rudy et al., (2017)) instead of the LASSO for ${\cal A}$. E-SINDy aggregates the coefficients $\hat{\beta}^{(b)}$ by selecting covariates whose inclusion probability is greater than some prespecified value $\tau\in(0,1)$. (We denote this selection procedure as ipτ in our experiments.) Of the variables with an inclusion probability greater than $\tau$, the $\hat{\beta}$ are set to the average fitted values across bags. In our experiments, we refer to this ensemble approach as $\tilde{\mathcal{A}}_{K,B}$.
5 Experiments
Our experiments assess the stability of five different variable selection approaches, described in Table 1, in two experimental settings.
NOTATION SELECTION CRITERION
We perform experiments in two settings: (a) identifying important variables in a synthetic dataset that contains a small collection of highly correlated coefficients, of which only a few are actually used to construct the response $y$ (§ 4.1) and (b) identifying the latent Lotka-Volterra governing differential equations from noisy trajectory data generated from these equations (§ 4.2). For each experimental setting, we generate independent datasets for $N$ trials, each contaminated with independent Gaussian noise, allowing us to examine the stability across a collection of trials. See supplement for details.
In each experiment, we compute sparse estimates of $\beta$ using both a base algorithm $\mathcal{A}$ and a subbagged version of $\mathcal{A}$ denoted $\tilde{\mathcal{A}}_{K,B}$. Based on (5), we compute the stability for trial $j\in[N]$ using
This definition of stability measures, for trial $j$, the proportion of LOO selected models $\hat{M}_{j}^{\setminus i}$ for $i\in[n]$ that have no overlap with the set of selected models $\hat{M}_{j}$ returned by training on the entire dataset. The left column of Figure 2 plots the empirical cumulative distribution function (CDF)
$\frac{1}{N}\sum_{j=1}^{N}\mathbbm{1}\{\delta_{j}\leq\delta\},$ (15)
as a function of $\delta\in[0,1]$ to highlight variation across trials with different random seeds. Curves higher on the plot are better in the sense that the corresponding method achieves model selection stability $\delta$ (Definition 2.2) for a larger fraction of trials, and our primary interest is in small $\delta$ (i.e., high stability).
$\frac{1}{N}\sum_{j=1}^{N}\frac{\mathbbm{1}\{m^{*}\in\hat{M}_{j}\}}{|\hat{M}_{j}|},$ (16)
where $m^{*}$ is the “true” data generating model. This measure down-weights the accuracy by the number of models returned $|\hat{M}_{j}|$. This metric is used to compare methods in the center column of Figure 2.
5.1 Penalized regression with LASSO
@@
Figure 1: Covariance matrix for each simulated dataset. Covariate indices 0 and 2 were used to construct the response $y$.
(a) Regression results. In the left plot, we plot the stability curves for the hyperparameters $\tau=0.3$, $k=2,$ and $\varepsilon=0.8$ since these values yield a utility-weighted accuracy of approximately 0.18 for each $\mathcal{S}$.
(b) Lotka-Volterra results. In the left plot, we plot the stability curves for the hyperparameters $\tau=0.63$, $k=2,$ and $\varepsilon=0.09$ since these values yield a utility-weighted accuracy of approximately 0.3 for each $\mathcal{S}$.
@@
Figure 2: (Left column) Empirical CDF for the stability measures $\delta_{j}$, computed in (14), across $j\in[N]$ for each $\mathcal{S}$. (Center column) Utility-weighted accuracy, computed in (16), versus the worst-case instability across $\delta_{j}$. (Right column) Median number of models returned across $j$ versus the maximum $\delta_{j}$ across $j\in[N]$ for each $\mathcal{S}$. (Center and right columns) We plot the values $k\in[1,\dots,6]$, and $\epsilon\in(0,1)$ and $\tau\in(0,1)$ with approximately evenly spread values across their supports.
Our first experiment is based on the motivating example in (9), where we have two important covariates and a small collection of highly correlated covariates. Figure 1 visualizes the covariance structure of the $20$ variables in the design matrix $X$, where the nonzero covariance between distinct covariates is 0.99. For each of the $N=200$ trials, we generate a dataset with $d=20$ covariates and $n=30$ training samples.
@@
Our learning algorithm $\mathcal{A}$ is the LASSO, which corresponds to solving (7) with $R(\beta)=||\beta||_{1}$, and $\tilde{\mathcal{A}}_{K,B}$ corresponds to a subbagged LASSO with $B=10,000$ bags and $K=25$ training samples per bag. We provide additional implementation details about this experiment in the supplement.
@@
Figure 2(a) provides an empirical CDF of $\delta$ (15), the utility-weighted accuracy (16) versus the empirical worst-case instability, and the median number of sets returned across $N$ trials versus the worse-case instability. The empirical CDF across our chosen selection methods $\mathcal{S}$ show that the selection methods that return a single model, the argmax and ipτ, are the least stable. Subagging, which helps in stabilizing the estimated weights $\hat{w}$ assigned to each model $m$, yields marginal gains in stability when using the argmax. Top-$k$ and the inflated argmax, on the other hand, provide enhanced stability due to the fact that these methods are able to return more than one model.
By plotting the utility-weighted accuracy compared to the worst-case instability, we hope to see a method that provides high utility-weighted accuracy while still providing stable model selections. The ipτ selection method fails to control the worst-case instability across a variety values of $\tau$, but is able to achieve good utility-weighted accuracy. Top-$k$ is able to achieve good utility-weighted accuracy with stable estimate. However, the inflated argmax is able to attain similar stability as top-$k$, while achieving a higher utility-weighted accuracy across a range of stability levels. This boost in utility-weighted accuracy is due to the adaptivity of the inflated argmax: this method will output a single model as often as possible, whereas top-$k$ will always output $k$ model selections, leading to a comparatively lower utility in many cases.
@@
5.2 Lotka-Volterra experiment
$\dot{u}^{(1)}(t)$ $=\;\;\,\alpha u^{(1)}(t)-\beta u^{(1)}(t)u^{(2)}(t),$ (17a)
$\dot{u}^{(2)}(t)$ $=-\gamma u^{(2)}(t)+\zeta u^{(1)}(t)u^{(2)}(t),$ (17b)
where $u^{(1)}$ is the population density of prey, $u^{(2)}$ is the population density of predators, $\dot{u}^{(1)}$ and $\dot{u}^{(2)}$ are the population growth rates, $t$ corresponds to time. The parameters $\alpha,\beta,\gamma,\zeta>0$ respectively refer to the prey’s growth rate, the predator’s impact on the prey’s death rate, the predator’s death rate, and the prey’s impact on the predator’s growth rate (Lotka,, 1925).
We generated synthetic datasets containing $n=21$ time points of abundance data for predator and prey populations, solved from the Lotka-Volterra equations and contaminated with Gaussian noise. $N=100$ trials were independently generated according to the same data generation process. We provide further details of this data generation process in the supplement. We visualize one set of trajectories in Figure 4.
We now imagine we do not know our system is governed by the Lotka-Volterra system and wish to discover the governing model from data using SINDy. To construct our library of functions, we selected degree-two polynomials as our covariates, and for one dataset, we visualize the corresponding covariate correlations in Figure 4. As seen in the correlation matrix, the covariate structure contains highly correlated variables (e.g., $u^{(1)}$ and $(u^{(2)})^{2}$), which impacts the model selection stability of SINDy.
In our experiments, we solve the optimization in (7) with the STRidge penalization (Rudy et al.,, 2017) as our base algorithm $\mathcal{A}$, as in Fasel et al., (2022).
Figure 2(b) shows that (a) selection methods that return a single model (argmax and ipτ, even with subagging the base algorithm) are less stable than methods that return at least one selected model (the inflated argmax and top-$k$), (b) the inflated argmax provides higher utility-weighed accuracy scores for similar worst-case stability levels as top-$k$, and (c) the inflated argmax provides smaller sets of returned models compared to top-$k$ for a wide range of worst-case stability values.
6 Discussion
The combination of bagging and the inflated argmax described in this paper can be applied to any (potentially unstable) black-box model selection procedure and offer stability guarantees described in Theorem 1. We define a notion of stability that is distinct from other works, one that computes the proportion of LOO selected models are disjoint from the selection made by training on the full data. This is in contrast to, for example, covariate-level stability in linear models.
Our experiments illustrate that, in addition to providing theoretical stability guarantees, our method outperforms ad hoc procedures. Specifically, choosing the most frequently selected model across bags provides limited stability gains relative to an unstable base algorithm. Selecting the top-$k$ most frequently selected models across bags enhances stability but is not adaptive to the underlying empirical uncertainty of the model selection process; $k$ models will be returned to the user regardless of the data. Choosing the returned model based on marginal inclusion probabilities of individual covariates is limited to covariate stability tasks and presents hyperparameter selection challenges. In contrast, the inflated argmax adapts to the number of sets returned based on model selection uncertainty, provides an approach for achieving a desired level of stability, and provides competitive utility-weighted accuracies.
Our framework provides multiple opportunities for further investigation, including exploring the interplay of stable model selection and identifiability and opportunities for the stabilization of black-box model selection methods without bagging (which may present computational barriers in some settings).
Acknowledgements
(a) $\varepsilon=0.00$
(b) $\varepsilon=0.15$
(c) $\varepsilon=0.30$
Figure 5: Visualization of the inflated argmax for increasing values of $\varepsilon$. At each corner of the triangle, $m_{1}$, $m_{2}$, and $m_{3}$ are three candidate models to select. When $\varepsilon=0$, this selection process is the same as the standard argmax. If a weight vector $\hat{w}$ lands in one of the colored regions, the inflated argmax returns the associated $\hat{M}$. For example, if $\hat{w}$ lands of the region where red and yellow overlap, models $\{m_{1},m_{2}\}$ are returned.
Appendix A Visualizing the inflated argmax
Figure 5 visualizes the model selection landscape for 3 candidate models for various choices of $\varepsilon$. As $\varepsilon$ increases, the inflated argmax is more likely to return more than one model, evidenced by the larger overlapping areas. Similarly, when $\varepsilon$ decreases, the inflated argmax becomes less and less likely to return more than one set, and in the limit when $\varepsilon\to 0$, the inflated argmax approaches the behavior of the standard argmax.
Appendix B Experiment details

### v2

@@ 5.1.1 Base algorithm $\mathcal{A}$: Sparse Identification of Nonlinear Dynamical Systems (SINDy)
5.1.2 Data generation
We generate synthetic datasets containing $n=21$ time points of abundance data for predator and prey populations, solve from the Lotka-Volterra equations and contaminate with Gaussian noise. This data is generated to match key characteristics of a real world predator-prey dataset from the Hudson Bay Company (Hewitt,, 1921) in terms of the periodicity of observations in relation to the dynamics. $N=100$ trials (i.e., independent datasets) are independently generated according to the same data generation process. We provide further details of this data generation process in § D.1. We visualize one set of trajectories in Figure 2.
To construct our library of functions, we selected degree-two polynomials as our covariates, and for one dataset, we visualize the corresponding covariate correlations in Figure 2. As seen in the correlation matrix, the covariate structure contains highly correlated variables (e.g., $u^{(1)}$ and $(u^{(1)})^{2}$), which impacts the model selection stability of SINDy.
In our experiments, we solve the linear regression problem with the STRidge penalization (Rudy et al.,, 2017) as our base algorithm $\mathcal{A}$, as in Fasel et al., (2022).
5.1.3 Results
We see that across various choices of $\varepsilon$, the inflated argmax provides enhanced stability while at the same time yielding small and accurate sets of selected model(s). This boost in utility-weighted accuracy is due to the adaptivity of the inflated argmax: this method will output a single model as often as possible, whereas for example, top-$k$ will always output $k$ model selections, leading to a comparatively lower utility in many cases. This point is further confirmed in the plot of the median number of models returned vs. worst-case instability in Figure 3.
For a fixed level of worst-case instability empirically seen, we can visualize the empirical CDF to assess the distribution of instability values $\delta_{j}$ across trials $j\in[N]$. We see in Figure 3 that the inflated argmax provides a distribution of instability values much closer to 0 (i.e., higher stability) compared to all other conventional approaches investigated.
In summary, Figure 3 shows that (a) selection methods that return a single model (argmax and ipτ, even with subagging the base algorithm) are less stable than methods that return one or more selected models (i.e., the inflated argmax and top-$k$), (b) the inflated argmax provides higher utility-weighed accuracy scores for similar worst-case stability levels as top-$k$, meaning that the inflated argmax is returning smaller sets containig the correct model more often and (c) the inflated argmax provides smaller sets of returned models compared to top-$k$ for a wide range of worst-case stability values.
Figure 3: Lotka-Volterra results. (Left column) Empirical CDF for the stability measures $\delta_{j}$, computed in (5), across $j\in[N]$ for each $\mathcal{S}$. We chose the parameters $\tau=0.63$, $k=2,$ and $\varepsilon=0.09$ since these values yield a utility-weighted accuracy of approximately 0.3 for each $\mathcal{S}$. (Center column) Utility-weighted accuracy, computed in (7), versus the worst-case instability across $\delta_{j}$. (Right column) Median number of models returned across $j$ versus the maximum $\delta_{j}$ across $j\in[N]$ for each $\mathcal{S}$. (Center and right columns) We plot the range of values $k\in[1,\dots,6]$, and $\epsilon\in(0,1)$ and $\tau\in(0,1)$ with approximately evenly spread values across their supports.
Figure 4: Empirical correlation matrix of the 11 proteins. Data from Sachs et al., (2005).
5.2 Graph selection on flow cytometry data
We consider a model selection problem where the model to select is the sparse connection structure within a graph. The base algorithm $\mathcal{A}$ we consider for this problem is the Graphical LASSO (gLASSO), which is developed in Friedman et al., (2008). The gLASSO estimates the structure of a graph by performing a maximum likelihood estimation of the precision matrix $\Theta$ subject to an $l_{1}$ penalty of the estimated precision matrix. More precisely, the gLASSO assumes the data $x_{i}\sim\mathcal{N}(0,\Theta^{-1})$, where $x_{i}\in\mathbb{R}^{d}$ across samples $i\in[n]$, and finds an estimate $\hat{\Theta}$ such that
$\hat{\Theta}=\mathop{\!\arg\!\max}_{\Theta}\ \log\det\Theta-\text{tr}(\tilde{\Sigma}\Theta)-\lambda||\Theta||_{1},$ (11)
where $||\cdot||_{1}$ is the entry-wise $l_{1}$ norm, tr()($\cdot$)( ⋅ ) is the trace, $\tilde{\Sigma}$ is the sample covariance matrix, and $\lambda$ is the penalization hyperparameter. In our experiments, we select $\lambda$ via cross-validation, as described in §D.2.1, and hold this quantity fixed across our experiments.
We compute LOO stability results from a flow cytometry dataset in Sachs et al., (2005) (labeled “cd3cd28icam2+u0126”), containing $d=11$ proteins (nodes) and $n=759$ cells (samples). Figure 4 shows the empirical correlation structure of these 11 proteins in this dataset, which clearly shows grouped relationships among proteins.
SELECTION LOO AVG. LOO
METHOD INSTABILITY SET SIZE
$\mathop{\!\arg\!\max}\circ\ \mathcal{A}$ 0.453 1.00
ip0.5 $\circ$ $\tilde{\mathcal{A}}_{K,B}$ 0.013 1.00
$\mathop{\!\arg\!\max}\circ\ \tilde{\mathcal{A}}_{K,B}$ 0.112 1.00
top-2 $\circ\ \tilde{\mathcal{A}}_{K,B}$ 0.008 2.00
$\ \mathop{\!\arg\!\max}^{0.02}$ $\circ\ \tilde{\mathcal{A}}_{K,B}$ 0.008 1.58
Table 2: Table of stability and the average number of model structures returned across LOO trials for selecting the structure of the inverse precision matrix for a flow cytometry dataset taken from Sachs et al., (2005). In this experiment, $\mathcal{A}$ is the graphical LASSO (Friedman et al.,, 2008), $\tilde{A}_{K,B}$ is the bootstrapped graphical LASSO with $B=10,000$ and $K=700$.
Figure 5: Visualization of the top two graph structures selected via subbagged gLASSO with 10,000 bags. The red connection shown in the right graph highlights that this connection is the only difference between the top 1 and top 2 selected graph structures. The top 1 graph was selected for 9.31% of bags, and the top 2 graph was selected for 8.54% of bags.
5.2.1 Results
Since we have one dataset to compute the LOO stability, $N=1$. We compute stability results for the algorithms gLASSO ($\mathcal{A}$) and subbagged gLASSO ($\tilde{\mathcal{A}}_{K,B}$), where $K=700$ and $B=10,000$, for the selection methods argmax, ip0.5, top-2, and argmax0.02. The value for $\varepsilon$ for the inflated argmax was chosen so that the LOO instability closely matched the LOO instability of top-2, and $\tau$ in ipτ was chosen so that the number of selected connections was approximately the the same as the number of connections given by the top selected subbagged graph. These results are shown in Table 2.
As seen in Table 2, our method, argmax0.02, provides the lowest LOO instability while minimizing the average number of graphs selected across the 759 LOO model fits. We also note that the percentage of model fits on bags that resulted in the top 2 models shown in 5 are relatively close: 9.31% for the top graph structure versus 8.54% for the second most selected graph structure, which suggests a reasonable degree of uncertainty in the “best” model. Due to this uncertainty in the selected model, selection methods that return only one model (i.e., argmax and ipτ) cannot sufficiently account for this uncertainty in the top model. Since the top two graph structures differ by one connection, this kind of result can provide evidence for follow up experiments, for example, to elucidate the relationship between the pakts473 and pmek proteins.
6 Discussion
The combination of bagging and the inflated argmax described in this paper can be applied to any black-box model selection procedure and offer stability guarantees described in Theorem 1. We define a notion of stability that is distinct from other works, one that computes the proportion of LOO selected models that are disjoint from the selection made by training on the full data. This is in contrast to, for example, covariate-level stability in linear models.
[…]
Acknowledgements
Appendix A Example: Linear regression with correlated covariates
@@
Appendix B Variable selection in linear models with LASSO
In the linear case, the data is modeled as the form
$\hat{m}(\mathcal{D})=\mathbbm{1}\{|\hat{\beta}|>0\}$ (18)
where $\hat{\beta}$ implicitly depends on the dataset $\mathcal{D}$. In this context, the LASSO provides unstable model selections, particularly in settings with nontrivial correlations among covariates (Meinshausen and Bühlmann,, 2010). Moreover, since many variable selection algorithms, including the LASSO, only return a single set of selected covariates, these methods do not directly provide an approach for assessing uncertainty in the selected model.
Appendix C Experiment: Penalized linear regression experiment with highly correlated covariates
@@
Figure 6: Covariance matrix for each simulated dataset. Covariate indices 1 and 3 were used to construct the response $y$.
@@
Figure 7: Regression results. In the left plot, we plot the stability curves for the parameters $\tau=0.3$, $k=2,$ and $\varepsilon=0.8$ since these values yield a utility-weighted accuracy of approximately 0.18 for each $\mathcal{S}$. (Left column) Empirical CDF for the stability measures $\delta_{j}$, computed in (5), across $j\in[N]$ for each $\mathcal{S}$. (Center column) Utility-weighted accuracy, computed in (7), versus the worst-case instability across $\delta_{j}$. (Right column) Median number of models returned across $j$ versus the maximum $\delta_{j}$ across $j\in[N]$ for each $\mathcal{S}$. (Center and right columns) We plot the values $k\in[1,\dots,6]$, and $\epsilon\in(0,1)$ and $\tau\in(0,1)$ with approximately evenly spread values across their supports.
This experiment is based on the motivating example in (12), where we have two important covariates and a small collection of highly correlated covariates.
C.1 Data generation
$x_{i}\sim\mathcal{N}(0,C),$ (19)
where $x_{i}\in\mathbb{R}^{d}$, $\mathcal{N}$ denotes a Normal distribution, and $C\in\mathbb{R}^{d\times d}$ is visualized in Figure 6. Figure 6 visualizes the covariance structure of the $20$ covariates in the design matrix $X$, where the nonzero covariance between distinct covariates is 0.99, corresponding to extremely correlated covariates. For a single trial $j$, we simulate a dataset with $d=20$ covariates and $i\in[n]$ samples. We take the first and third columns as the covariates used to create the response $y$,
C.1.1 Optimization hyperparameter
C.2 Small sample size ($n=30$)
We generate independent datasets (i.e., trials) for a small sample size, $n=30$. For each of the $N=200$ trials, we generate a dataset with $d=20$ covariates and $n=30$ training samples.
C.2.1 Base algorithm $\mathcal{A}$: LASSO
@@
Our learning algorithm $\mathcal{A}$ is the LASSO, which corresponds to solving (17) with $R(\beta)=||\beta||_{1}$, and $\tilde{\mathcal{A}}_{K,B}$ corresponds to a subbagged LASSO with $B=10,000$ bags and $K=25$ training samples per bag.
C.2.2 Results
@@
Figure 7 provides an empirical CDF of $\delta$ (6), the utility-weighted accuracy (7) versus the empirical worst-case instability, and the median number of sets returned across $N$ trials versus the worse-case instability. The empirical CDF across our chosen selection methods $\mathcal{S}$ show that the selection methods that return a single model, the argmax and ipτ, are the least stable. Subagging, which helps in stabilizing the estimated weights $\hat{w}$ assigned to each model $m$, yields marginal gains in stability when using the argmax. Top-$k$ and the inflated argmax, on the other hand, provide enhanced stability due to the fact that these methods are able to return more than one model. The ipτ selection method fails to control the worst-case instability across a variety values of $\tau$, but is able to achieve good utility-weighted accuracy. Top-$k$ is able to achieve good utility-weighted accuracy with stable estimates. However, the inflated argmax is able to attain similar stability as top-$k$, while achieving a higher utility-weighted accuracy across a range of stability levels.
@@
C.3 Larger sample size ($n=300$)
[…]
$\left\{\{x^{(1)},x^{(2)}\}\times\{x^{(3)},x^{(4)},x^{(5)}\}\right\}.$
Appendix D Main experimental details